OpenStudy (anonymous):

A car dealership interviewed every driver who came onto their lot on Tuesday. This produced the following results: there were 61 people who owned trucks and 115 people who owned cars. If 14 people owned both a truck and a car and 8 people owned neither, how many people were interviewed

OpenStudy (anonymous):

184

OpenStudy (anonymous):

61+115=176 14 own cars and trucks, this means they have been counted twice so: 176-14 = 162 and 8 people own neither so haven't been counted yet, so: 162 +8 = 170